OpenStudy (anonymous):

what is 19/20 in a percentage

OpenStudy (ccswims):

Just do what everyone said in the last post

OpenStudy (ccswims):

19 divided by 20 times 100

OpenStudy (lexy148):

its 95% (use a calculator)

TheSmartOne (thesmartone):

\[\frac{ 19 }{ 20 }\times \frac{ 5 }{ 5}= \frac{ 95 }{100 }= 95%\]

TheSmartOne (thesmartone):

Hmm it didn't put the % sign after 95...

OpenStudy (xxdragoonxx):

You could reduce it to 95/100=0.95*100=95/5=19/20